I would appreciate some help. It has been a while since I have done this. I am to solve the equation: 2/5 + 1/3 t = 5.
Thank you very much.

Answer by sdmmadam@yahoo.com(530)

Is the problem (2/5) + (1/3) t = 5 ?
2/5 + (1/3) t = 5. ----(1)
(1/3)t = 5 -2/5 (subtracting 2/5 from both the sides or shifting 2/5 from the left to the right side to group like terms, changing sign while changing side)
t/3 =(25-2)/5
t/3 = 23/5
Multipllying by 3
t = (23/5)X3 = 69/5
Answer: t = 69/5
Verification: Putting t =69/5 in (1)
LHS = (2/5)+(1/3)X(69/5)
=2/5 +23/5
=(2+23)/5
=25/5
=5 = RHS
Therefore our value is right.
Note: If the given problem is 2/5 + 1/(3t) = 5. ----(2)
Then 1/(3t) = 5-2/5
1/3t = 23/5
(3t)X(23) = (1)X(5) (cross multiplying)
[the nr of the left multiplied by the dr of the right = the nr of the right multiplied by the dr of the left]
69t = 5
Therefore t = 5/69
Verification:Putting t =5/69 in (2)
LHS = (2/5)+ [1/(3X5/69)]
2/5+23/5 = 25/5 =5 =RHS
Therefore our value is right